OpenStudy (anonymous):

What are the factors of x2 + 12x + 20? (x + 4)(x + 5) (x + 2)(x + 10) (x − 4)(x − 5) (x − 2)(x − 10)

OpenStudy (anonymous):

1.Since the constant is positive, its factor will either be both positive or both negative,. 2.Think of the factors of 20 that when added up will equal to 12. 3. You can go from there

Directrix (directrix):

x2 + 12x + 20 = x² + 10x + 2x + 20 = x(x+ 10) + 2(x + 10)= (x + 10) (x + 2). @harliedanielfrye

Directrix (directrix):

>>2.Think of the factors of 20 that when added up will equal to 12. Those are 10 and 2. The 12x in the original problem was split into 10x + 2x in preparation for the factoring by grouping technique.